# LeanCat: A Benchmark Suite for Formal Category Theory in Lean (Part I: 1-Categories)

### v1

@@ 3.1 Evaluation Protocol
Verification: A proof attempt is considered successful if the Lean theorem prover accepts it as a valid proof of the given statement. We instrument Lean to check the model outputs automatically. If the model output is incomplete or incorrect (does not typecheck), that attempt counts as a failure. Models do not get to “see” the result of the verification; in pass@k evaluation, each attempt is separate.
Pass@k Calculation: We compute pass@1 as the fraction of problems for which the model produced a correct proof in a single attempt. Pass@4 is the fraction of problems where at least one out of 4 attempts succeeded. With 100 problems, these percentages translate directly to number of problems solved. We note that LeanCat’s problems are all of roughly equal weight (each is a standalone theorem), so simple pass rates are a meaningful indicator of overall capability. We also track pass@1 within each difficulty category (Easy/Med/High) to see how performance degrades as difficulty increases.
We used a uniform evaluation setup: each attempt had an output budget of 50,000 tokens and a Lean verification time limit of 5 minutes, and all models ran against the same Lean environment (Lean 4 + Mathlib 4.19.0) to ensure consistency. If a model exceeded the token budget or failed to finish verifying within the time limit, the attempt was counted as a failure. In practice, however, these resource limits were rarely the deciding factor: most attempts either found a proof quickly (typically within 30 seconds, except for reasoning model like DeepSeek) or got stuck almost immediately (often after only a few dozen tokens).
We emphasize that pass@4 is not meant as a realistic usage scenario (where one would run a model four times per theorem); rather, it provides an optimistic upper bound on success if we could perfectly pick from a handful of model tries. In an ideal case where attempts are independent, pass@4 could be significantly higher than pass@1, but as we will see, the increase was modest for LeanCat. This suggests that when a model fails on a problem in one attempt, trying again usually yields a similar outcome unless guided by a different strategy.
Table 1: LeanCat FL Solved by LLM (Pass@1/Pass@4)
Model Easy Medium High All
claude-opus-4-5 32.50% / 50.00% 4.17% / 4.76% 0.00% / 0.00% 8.25% / 12.00%
gpt-5.2 27.50% / 30.00% 0.00% / 0.00% 0.00% / 0.00% 5.50% / 7.00%
gemini-3-pro 11.25% / 25.00% 2.38% / 7.14% 0.00% / 0.00% 3.25% / 8.00%
deepseek-reasoner 18.75% / 40.00% 0.60% / 2.38% 0.00% / 0.00% 4.00% / 9.00%
kimi-k2-0905 10.00% / 20.00% 0.00% / 0.00% 0.00% / 0.00% 2.00% / 4.00%
Preliminary data indicates only a 1-2 problem increase from pass@1 to pass@4 for the top models, reinforcing the difficulty of LeanCat tasks.
LeanBridge: LeanBridge implements a retrieve–analyze–generate–verify loop that augments an LLM with mathlib retrieval and verifier feedback. Given a problem, we first use the natural-language statement as a query to LeanExplore to retrieve relevant mathlib entities (e.g., definitions, lemmas). The retrieved snippets are then provided to the model as context for analysis and Lean code generation. Each generated proof script is checked in a clean Lean environment; a candidate is accepted only if it typechecks and contains no sorry or admit. To guard against superficial “green” proofs that do not match the intended meaning, all accepted candidates are additionally reviewed by a human expert for semantic correctness with respect to the original problem statement. If verification fails, LeanBridge parses the compiler error messages to decide whether additional retrieval is needed; it then augments the context with both the retrieved information and the verifier feedback, and prompts the model to revise the proof. Unless otherwise stated, we cap the loop at a maximum of 4 iterations for both (i) natural-language-to-statement formalization and (ii) natural-language theorem proving.
3.2 Baseline Results and Analysis
We evaluated five state-of-the-art models on LeanCat under the protocol described above. We summarize the main findings below.
  • Overall success remains low. On pass@1 (first attempt), the best model (Claude Opus 4.5) solved $8.25\%$ problems; GPT-5.2 solved $5.5\%$, DeepSeek Reasoner solved $4\%$, while Gemini 3 Pro at $3.25\%$ and Kimi at $2\%$. Across all models, only 10 distinct problems were solved by at least one model on the first attempt, leaving $91.75\%$ unsolved at pass@1. Allowing up to four attempts per problem improves results but does not change the overall picture: pass@4 rises to $12\%$ for Claude Opus 4.5, $9\%$ for DeepSeek Reasoner and $8\%$ for Gemini 3 Pro, $7\%$ for GPT-5.2, and $4\%$ for Kimi. In total, 14 distinct problems are solved by at least one model at pass@4.
  • Clear Easy–Medium–High gap. Performance decreases monotonically with our difficulty annotations. For example, Claude Opus 4.5 achieves $32.5\%$ on Easy, $4.17\%$ on Medium, and $0\%$ on High at pass@1 (and $50\%$, $4.76\%$, $0\%$ at pass@4). GPT-5.2 shows a similar trend ($27.5\%$, $\%, 00\%$ at pass@1). Even within the Easy subset, the absolute success rates remain far from saturation, suggesting that LeanCat’s “baseline difficulty” is already beyond what current models reliably handle once nontrivial abstraction and library navigation are required.
  • Case study (typical success): Problem 82. This problem is consistently solved by models that effectively translate the categorical notion of simplicity into concrete linear algebra. The successful approach recognizes that a simple object in $\mathrm{Vect}_{k}$ must be one-dimensional, then constructs an explicit isomorphism with k by leveraging a nonzero vector and the simplicity condition. The proof elegantly bridges abstract category theory with elementary vector space properties, demonstrating a clear understanding of how structural definitions manifest in concrete categories.
  • Retries help selectively, indicating high variance and brittle search. Moving from pass@1 to pass@4 yields only modest absolute gains for the strongest model (+5 for Claude Opus 4.5), but it substantially benefits some weaker models (e.g., DeepSeek Reasoner improves from 3 to 8). This pattern is consistent with high-variance behavior: many problems are either solved quickly or not approached effectively at all, and additional attempts mainly help when a model happens to sample a viable strategy or recall the right library lemma.
  • Error analysis: library knowledge gaps dominate, followed by abstraction mismatch and incomplete plans. Manual inspection of failed runs suggests three recurring failure modes: (i) library knowledge gaps: models often fail to recall the correct Mathlib definitions/lemmas or their usable forms, leading to dead ends or hallucinated lemma names; (ii) abstraction mismatch: some attempts switch to element-wise reasoning when the intended proof is categorical/structural, which is typically unproductive in Lean without substantial setup; (iii) incomplete multi-step planning: models may discharge a few local goals but then stall, unable to connect intermediate facts into a coherent end-to-end proof. Pure syntax-level errors occur, but are less frequent than these semantic/strategic failures.
Overall, these baselines confirm that LeanCat is substantially harder for current LLM-based provers than earlier Lean benchmarks. Even with multiple attempts, success remains sparse on Medium/High problems, pointing to a need for improved library retrieval, better abstraction-aware proof planning, and more reliable strategy exploration.
4 Discussion and Future Work
LeanCat as a benchmark (and a series). LeanCat is intended as a reusable checkpoint for LLM-based theorem proving in abstract mathematics. This paper introduces LeanCat-1 (1-category theory); we view it as the first installment of a broader LeanCat series, with planned extensions targeting richer categorical interfaces such as monoidal categories and higher-categorical structures (e.g., bicategories / strict 2-categories) that are already represented in Mathlib’s ecosystem.
Library integration. All LeanCat statements are formalized in Lean 4; as solutions are found, they can be merged into Mathlib, creating a feedback loop: benchmark $\rightarrow$ solutions $\rightarrow$ stronger libraries/solvers $\rightarrow$ harder remaining frontier.
What LeanCat stresses. Our results highlight three persistent bottlenecks for current automated provers: (i) library awareness (finding and applying the right Mathlib lemmas), (ii) abstraction control (staying at the correct categorical level rather than drifting to element-wise reasoning), and (iii) long-horizon coherence (maintaining a consistent plan across many dependent steps).
Future work and broader impact. On the benchmark side, we will expand LeanCat beyond 1-categories to include additional clusters and multi-theorem tasks, and to progressively cover higher abstraction layers—e.g., a monoidal track and a 2-categorical track (where monoidal categories can be viewed through the lens of a single-object bicategory), enabling finer-grained diagnosis of where provers fail as abstraction increases. On the solver side, promising directions include stronger retrieval over Mathlib, hierarchical decomposition into auxiliary lemmas, and multi-agent pipelines (planner/verifier/lemma-suggester). For human mathematics, we expect LeanCat-style checkpoints to help prioritize missing interfaces and reusable lemmas in the library; for AI, they provide a concrete target for improving abstraction-aware planning and library-grounded reasoning. Finally, porting LeanCat to other proof assistants (e.g., Coq or Isabelle) would enable cross-system comparisons and encourage transfer of proof-engineering techniques.
Acknowledgments
The Lean workshop at Westlake University was supported by the National Natural Science Foundation of China (NSFC), Grant No. 12574176. We thank Prof. Jian Li for supporting the organization of the workshop. We thank Xingyu Ren for inspiring this benchmark. We thank Prof. Lifan Guan, Huayi Chen, Bin Dong and Hao Zheng for helpful comments. We also thank Jiadong Adam Guo, Jinyuan Zhang, and Xiaodong Hu for providing API access for our large language model experiments.

### v2

@@ D.1 System Prompts
D.2 LeanBridge Prompts
For the LeanBridge agent, we use a distinct prompt structure to facilitate the retrieve-generate-verify loop:
Listing 2: Prompt template for LeanBridge proof generation
"""You are an expert in Lean 4 and Mathematics. Please finish the following proof in Lean4 code.
if provided a formal statement, do not change the original formal statement. Copy the final statement to prove exactly. Please include the complete header (including imports and namespaces) so that your code can pass the Lean4 compiler.
Please solve the statement step by step and provide your complete Lean4 code between ‘‘‘lean4 and ‘‘‘ after careful reasoning.
The statement for you to complete is:
Natural language statement:
{nl_statement}
Lean 4 formal statement:
{formal_statement}
Reference Knowledge (Search Results from Mathlib):
{search_contents}
Listing 3: Prompt template for LeanBridge refinement
"""The previous Lean 4 proof attempt contains errors. Your goal is to fix them according to the compilers error message.
Input:
Lean 4 code:
{original_proof}
Error messages:
{error_messages}
Instructions:
1. First, analyze the error message carefully in a "## Error Analysis" section. Identify specifically why the verification failed (e.g., missing import, wrong tactic, type mismatch, or a missing lemma).
2. Based on your analysis, decide whether you need to search Mathlib for external knowledge (definitions, theorems, or syntax) or if you can fix it immediately.
Condition A: If you need external information
Output the search query wrapped in the tag below. Make your query specific (e.g., a math statement, a definition name, or a type signature).
Format:
## Error Analysis
(Your analysis of the error and what you need to find)
[SEARCH: your_search_query]
Condition B: If you can fix it immediately
Output the analysis and the fully refined code.
Format:
## Error Analysis
(Your analysis of the error and the fix plan)
## Refined Code
‘‘‘lean
-- The complete, corrected Lean 4 code
D.3 Hyperparameters
Table 5: Detailed experimental configuration and hyperparameters.
Component Parameter Value / Description
LLM Generation Temperature $1.0$ (Encouraging diversity)
Max Tokens $50,000$ (Generalist models)
Model-specific max (Specialized provers)
Verification Lean Version v4.19.0
Timeout 300300 seconds per attempt
LeanBridge Loop Limit Max $4$ refinement iterations
Retrieval Top-33 results per search query
Appendix E Case Study
This section presents a detailed case study of a specific problem, including the original problem statement, the generated Lean 4 code, and the evaluation of the generated code.
E.1 Success Case
As depicted in Appendix Figure 3, Problem 59 is a representative success case: it is the only High-difficulty problem solved in our setting, achieved only by GPT-5.2 and Gemini-3-Pro with LeanBridge. We study the GPT-5.2 attempt. This run fails in early attempts but verifies successfully on the final refinement (the proving trajectory showed in Figure 8).
Problem 59 Proving Trajectory
 Refer to caption Refer to caption
Figure 8: Case Study (Success): Iterative Refinement on Problem 59. The proving trajectory by LeanBridge (GPT-5.2), showing the progression from initial attempts to a successful formal proof.
1. Strategy used (and mathematical correctness).
The final successful witness takes
$C:=\mathrm{Type}\,u_{1},\qquad D:=\text{a terminal category on }\texttt{PUnit},\qquad F:C\to D\text{ the constant functor}.$
This strategy is mathematically appropriate: limit-lifting becomes easy because $C$ has all limits and $D$ has trivial cone structure, while non-faithfulness is immediate because $F$ collapses distinct morphisms to the unique morphism in $D$.
2. Why it fails initially, and how LLM fixes it.
The first attempt tries to exploit “vacuity” by taking an empty codomain (e.g. PEmpty) and concludes lifting should be trivial. This fails due to (i) universe/typeclass issues when synthesizing the required Category instances and (ii) ill-typed constructions (e.g. attempting to inhabit PEmpty). In subsequent iterations, the run pivots to the clean terminal-codomain construction above, but still encounters Lean-engineering obstacles: universe mismatches when using Bool : Type 0 as an object of Type u_1, and brittle pattern matching tactics (e.g. casesm*). The final repair commits to the clean construction and patches these issues by using ULift Bool : Type u_1 (with explicit ULift.up) and replacing brittle matches with simple cases.
3. Why the fix works.
Limit lifting succeeds because Type u_1 already provides canonical limit data limit.cone K and limit.isLimit K, which can be used verbatim as the lift. The required cone isomorphism is trivial in a terminal category: all cone legs are definitional equal to the unique morphism, so the isomorphism can be constructed by extensionality with reflexive components. Non-faithfulness is witnessed concretely by two distinct endomorphisms on ULift Bool (identity vs. boolean negation), which both map to the unique morphism in $D$; this contradicts Functor.map_injective by evaluating at ULift.up true.
In summary, the run succeeds because it converges to a canonical, Lean-friendly counterexample: a constant functor from a complete category (Type) to a terminal category. This choice simultaneously trivializes the lifting obligation (via existing limit infrastructure in Type) and makes non-faithfulness explicit (via collapsing distinct functions), while the final patches resolve the remaining universe and pattern-matching pitfalls.
E.2 Failure Case
We also study the GPT-5.2 with LeanBridge. As depicted in Appendix Figure 4, Problem 63 is a representative failure case (the proving trajectory showed in Figure 9): combining both a mathematical-strategy error and a “lazy” failure mode.
Problem 63 Proving Trajectory
 Refer to caption Refer to caption
Figure 9: Case Study (Failure): Iterative Refinement on Problem 63. The proving trajectory by LeanBridge (GPT-5.2).
The task is to formalize in Lean + mathlib the claim that in $\mathcal{G}\mathrm{rp}$, categorical finite presentability matches finite group presentation (generators/relations), and that every group is a filtered colimit of finitely presented groups. In the file, “finitely presented as a group” is encoded by a custom predicate
$\texttt{IsFinitelyPresentedGrp }X$={} $\exists(\alpha,\mathrm{rels}),\texttt{Finite }\alpha\ \wedge$
$\texttt{rels.Finite }\wedge\texttt{Nonempty}(X\simeq_{*}\texttt{PresentedGroup rels})$
and the goals are (i) an equivalence
$\texttt{IsFinitelyPresentable (Grp.of X)}\;\leftrightarrow\;\texttt{IsFinitelyPresentedGrp X},$
and (ii) a filtered diagram $F$ landing in groups satisfying IsFinitelyPresentedGrp whose colimit is isomorphic to $X$.
1. Strategy used (and mathematical correctness).
The LLM aims to reuse existing mathlib results: reduce the statement to mathlib’s notion IsFinitelyPresentable in Grp, then connect that to the concrete PresentedGroup-based predicate, and finally invoke a filtered-colimit theorem in Grp. This high-level plan matches the correct mathematical strategy: one normally proves the equivalence via standard theorems (Baer/Freyd-style characterizations in algebraic categories) and then uses the classical fact that any group is a filtered colimit of finitely presented groups. However, when it cannot find the required bridge theorems, the LLM falls back to an invalid shortcut: it tries to realize the filtered colimit using a trivial/constant diagram (e.g. indexed by PUnit), which would effectively force IsFinitelyPresentedGrp X for arbitrary $X$, a statement that is false in general.
2. Why it fails, and how the LLM tries to fix it.
In the first attempt the LLM tries to close the equivalence using a guessed lemma name like Grp.isFinitelyPresentable_iff, but Lean reports it as an unknown constant (the lemma does not exist under that name). For the colimit part it attempts to define a constant functor into a filtered index category and then discharge the remaining obligations with automation (e.g. aesop), but the hard goals remain: one must actually prove IsFinitelyPresentedGrp (F.obj j) for each index $j$ and produce Nonempty (X $\simeq_{*}$ colimit F). After this failure the LLM “fixes” by switching placeholders: it inserts fail in place of the missing mathematics (which Lean rejects), then replaces fail by simp (which makes no progress on genuinely nontrivial goals), and finally tries contradiction-style discharge, attempting to reduce obligations to False and solve them via assumption or rcases. These last steps fail because no contradiction is derivable from the given hypotheses, and because some tactic steps are applied to ill-typed or non-inductive targets, indicating the repair loop has drifted away from the correct categorical structure.
3. Why the fixes do not work.
The fixes fail for structural reasons rather than minor syntax. First, the proof fundamentally needs nontrivial “bridge” theorems (or a substantial new development) connecting mathlib’s categorical finite presentability in Grp with the specific user-defined predicate IsFinitelyPresentedGrp based on PresentedGroup. The LLM never finds or constructs this bridge; instead it guesses lemma names. Second, the constant-diagram shortcut cannot satisfy the statement “every group is a filtered colimit of finitely presented groups,” since it would imply every group is itself finitely presented. Third, swapping in fail, simp, or attempting to conjure False does not create the missing mathematical content; Lean correctly rejects these as non-proofs.
In summary, the run fails because the LLM does not supply the missing mathematical infrastructure required by the formal statement: (i) a correct, explicit connection between IsFinitelyPresentable in Grp and the concrete PresentedGroup-based predicate, and (ii) a correct nontrivial filtered diagram witnessing $X$ as a filtered colimit of finitely presented groups. When those are not available, the LLM falls back to lemma-name guessing and tactic placeholders, which do not typecheck or do not solve the goals, so verification never succeeds.
Appendix F Datasheet for Datasets
  • Motivation: To benchmark library-level abstraction reasoning in LLMs.
  • Composition: 100 pairs of Natural Language and Lean 4 statement code.
  • Collection Process: Expert curation from textbooks, unpublished lecture note and research papers.
  • Maintenance: Hosted on GitHub under MIT.